OpenStudy (anonymous):

solve ln(x)+ln(2x)=2

OpenStudy (australopithecus):

apply this rule ln(x) + ln(y) = ln(y*x)

OpenStudy (australopithecus):

and combine the ln expressions

OpenStudy (australopithecus):

then make both sides exponents to the base e as e^(ln(x)) = x

OpenStudy (anonymous):

write it as one natural log: ln (2x^2) = 2 write it exponential form: e^2 = 2x^2 divide by 2: e^2/2 = x^2 square root both sides: plus/minus (1/sqrt2) * e = x

OpenStudy (anonymous):

then check... the solution... one does not work...

OpenStudy (anonymous):

this is still confusing. Is there an easier way

OpenStudy (anonymous):

both our solutions utilized the properties of logs: \[\large \log_b {MN} = \log_bM + \log_bN\] \[\large \log_b {\frac {M}{N}} = \log_bM - \log_bN\] \[\large \log_bM^n=n*\log_bM\]

OpenStudy (anonymous):

you'll just have to just practice and get used to using these properties..

OpenStudy (anonymous):

well, I have to know this in an hour.